OpenStudy (anonymous):

what is the vertex of y= -2x2 + x +1

OpenStudy (anonymous):

Remember how you found b earlier?

OpenStudy (anonymous):

foil?

OpenStudy (anonymous):

Here, a=-2, b=1, c=1

OpenStudy (anonymous):

Yes, but you don't have to foil here, just recognize a,b,c

OpenStudy (anonymous):

The vertex of a parabola written in this form ax^2+bx+c is the point \[(\frac{-b}{2a},f(\frac{-b}{2a}))\]

OpenStudy (anonymous):

So you can find the x-coordinate by evaluation -b/(2a)

OpenStudy (anonymous):

so 1/a(-2)

OpenStudy (anonymous):

1/2(-2)

OpenStudy (anonymous):

remember the negative sign

OpenStudy (anonymous):

for b

OpenStudy (anonymous):

-b/(2a)=-1/(2*(-2))

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

but you're not done yet

OpenStudy (anonymous):

you have the x-coordinate, which is 1/4, what is the y-coordinate?

OpenStudy (anonymous):

to get the y-coordinate plug in (1/4) for x in the equation

OpenStudy (anonymous):

okay